# Reliable algorithm selection for machine learning-guided design

### v1

@@ B.1 More general success criteria
The main text considers success criteria of the following form: $\theta_{\lambda}=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]\geq\tau$ for some $g:\mathcal{Y}\rightarrow\mathbb{R},\tau\in\mathbb{R}$. More generally, we can use prediction-powered inference techniques to compute valid p-values whenever $\theta_{\lambda}$ can be expressed as the minimizer of the expectation of some convex loss [2]:
$\theta_{\lambda}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[\ell_{\theta}(X,Y)],$
where $\ell_{\theta}$ is convex in $\theta$. For example, when $\ell_{\theta}(X,Y)=(g(Y)-\theta)^{2}$ for some $g:\mathcal{Y}\rightarrow\mathbb{R}$, we recover the special case of $\theta_{\lambda}=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]$. We could not conceive of practical settings requiring the general characterization of $\theta_{\lambda}$, but it may be useful for future work.
@@ B.2 Design and labeled densities known up to normalizing constants
We can compute asymptotically valid p-values if we have unnormalized forms of the design and labeled densities, such as when sequences are generated from energy-based models [7], Potts models [34, 16], or other Markov random fields. Specifically, assume we can evaluate $p^{\text{u}}_{X;\lambda}(x)=a\cdot p_{X;\lambda}(x)$ and $p^{\text{u}}_{\text{lab}}(x)=b\cdot p_{\text{lab}}(x)$ for unknown constants $a,b\in\mathbb{R}$, where the superscript indicates that the densities are unnormalized. To leverage these in place of the exact densities in Algorithm 2, consider the corresponding scaled density ratios on the labeled data, $w^{\text{u}}_{j}:=p^{\text{u}}_{X;\lambda}(x_{j})/p^{\text{u}}_{\text{train}}(x_{j}),j=1,\ldots,n$. The self-normalized importance-weighted estimate of prediction bias,
@@
is a consistent estimator of the rectifier, $\mathbb{E}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[g(Y)-g(f(X))]$, just as $\hat{\Delta}$ is in Algorithm 2. Since $\hat{\Delta}^{\text{u}}$ is a ratio of means, the delta method can be used to derive its asymptotic distribution [29], whose variance can be estimated as
@@
We can then compute an asymptotically valid p-value as
$P:=1-\Phi\left(\left(\hat{\theta}-\tau\right)/\sqrt{\hat{\sigma}^{2}_{\hat{y}^{\prime}}/N+\hat{\sigma}^{2}_{\hat{\Delta}^{\text{u}}}}\right),$
where $\hat{\theta}$ is computed using $\hat{\Delta}^{\text{u}}$ instead of $\hat{\Delta}$.
@@ C.1 Prediction-only method
Inputs: Predictions for designs, $\{\hat{y}^{\prime}_{i}\}_{i=1}^{N}$; desired threshold, $\tau\in\mathbb{R}$.
@@
1: $\hat{\theta}\leftarrow\frac{1}{N}\sum_{i=1}^{N}g(\hat{y}^{\prime}_{i})$
2: $\hat{\sigma}^{2}_{\hat{y}^{\prime}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}(g(\hat{y}^{\prime}_{i})-\hat{\theta}))^{2}$
3: $P\leftarrow 1-\Phi\left((\hat{\theta}-\tau)/\sqrt{\hat{\sigma}^{2}_{\hat{y}^{\prime}}/N}\right)$
The prediction-only method runs multiple testing (Alg. 1) with p-values computed using only the predicted property values of the designs (Alg. 6), treating them as if they were labels. These p-values are asymptotically valid for testing whether $\mathbb{E}_{X\sim P_{X;\lambda}}[g(f(X))]\geq\tau$—that is, whether the expected function of predictions for designs, but not necessarily their labels, surpasses a threshold.
@@ C.2 Gaussian-mixture-model forecasts method
The GMMForecasts method follows Wheelock et al. [52], who model the forecast for a given sequence as mixture of two Gaussians (representing beliefs over the label if the sequence is “nonfunctional” and “functional”) with sequence-specific means, variances, and mixture proportion. To construct these forecasts, the method first assumes access to a predictive mean and variance for each sequence. When using predictive models that were ensembles (i.e., the feedforward ensemble in the GB1 experiments, and the feedforward and convolutional ensembles in the RNA experiments), we set these to the mean and variance, respectively, of the predictions for a sequence. When using the ridge regression model in the RNA experiments, the model weights were fit on $90\%$ of the training data, and the predictive mean for a sequence was set to the model’s prediction. The (homogeneous) predictive variance was set to the model’s mean squared error over the remaining $10\%$ of the training data. Wheelock et al. [52] then use the training data to fit a mapping from these initial predictive means and variances to the Gaussian mixture model (GMM) parameters. They also seek to address covariate shift between the design and training data by using a sequence’s edit distance from a reference training sequence (set to the wild type GB1 in the GB1 experiments, and the seed sequence in the RNA experiments) as an additional feature in fitting this mapping.
The forecasts also involve a key hyperparameter, $q\in[0,1]$. After the GMM parameters are inferred for each design, the method adjusts the sequence-specific mean of the “functional” Gaussian by taking a convex combination of it and the original predictive mean, where $q$ weights the latter. Using different values of $q$ reflects how much one trusts the initial predictions; high values result in forecasts where the “functional” Gaussian mean is determined largely by the original predictive mean. We ran the method with $q\in\{0,0.5,1\}$ to span the range of possible values.
@@ C.3 Calibrated forecasts method
The CalibratedForecasts method invokes the idea that the forecasts for sequences, or models of the conditional distributions of the label, should be statistically consistent with the true conditional distributions of the label. Specifically, we want the forecasts to be calibrated as defined by Kuleshov et al. [26], whose definition is related to the notion of probabilistic calibration [17]:
@@
where $F_{X}$ denotes the CDF of the forecast for a sequence $X\in\mathcal{X}$, and the probability is over the distribution of labeled data. That is, for any $p\in[0,1]$, the label will fall under the $p$-quantile given by a calibrated forecast with frequency $p$.
For any given sequence, let an initial forecast be a Gaussian with mean and variance set to a predictive mean and variance, as described above for GMMForecasts . We then use the held-out labeled data to learn a transformation of these initial forecast CDFs, with the objective of achieving calibration (Eq. 12). Specifically, we use isotonic regression, following Kuleshov et al. [26]. We then construct the forecast for any given design by first forming the initial forecast, and then transforming the corresponding CDF with the fitted isotonic regression function.
@@ C.4 Conformal prediction method
Inputs: Designs generated with each configuration, $\{x^{\lambda}_{i}\}_{i=1}^{N}$ for all $\lambda\in\Lambda$; predictive models used by each configuration, $\{f_{\lambda}\}_{\lambda\in\Lambda}$; held-out labeled data, $\{(x_{i},y_{i})\}_{i=1}^{n}$; desired threshold, $\tau\in\mathbb{R}$; error rate, $\alpha\in[0,1]$.
@@
2:      $\hat{y}^{\lambda}_{i}\leftarrow f_{\lambda}(x^{\lambda}_{i}),i=1,\ldots,N$triangleright predictions for designs
3:      $\hat{y}_{j}\leftarrow f_{\lambda}(x_{i}),j=1,\ldots,n$triangleright predictions for labeled data
4:      $v^{\lambda}_{i}\leftarrow\textsc{DensityRatio}_{\lambda}(x^{\lambda}_{i}),i=1,\ldots,N$triangleright density ratios for designs
5:      $w^{\lambda}_{j}\leftarrow\textsc{DensityRatio}_{\lambda}(x_{j}),j=1,\ldots,n$triangleright density ratios for labeled data
6:      $l^{\lambda}_{i}\leftarrow\textsc{SplitConformalLB}((\hat{y}^{\lambda}_{i},v^{\lambda}_{i}),\;\{(y_{j},\hat{y}^{\lambda}_{j},w^{\lambda}_{j})\}_{j=1}^{n},\;\alpha/(|\Lambda|\cdot N)),i=1,\ldots,N$triangleright Alg. 8
7:      $L_{\lambda}\leftarrow(1/N)\sum_{i=1}^{N}l^{\lambda}_{i}$triangleright average of designs’ conformal lower bounds
@@
Inputs: a design’s prediction and density ratio, $(\hat{y}^{\prime},w^{\prime})$; labels, predictions, and density ratios for held-out labeled data, $\{(y_{j},\hat{y}_{j},w_{j})\}_{j=1}^{n}$; $\alpha\in[0,1]$.
@@
1: $u_{j}\leftarrow\frac{w_{j}}{\sum_{j=1}^{n}w_{j}+w^{\prime}},j=1,\ldots,n$
2: $u^{\prime}\leftarrow\frac{w^{\prime}}{\sum_{j=1}^{n}w_{j}+w^{\prime}}$
3: $r\leftarrow\text{$(1-\alpha)$-quantile of the distribution comprising the mixture of point masses:}\sum_{j=1}^{n}u_{j}\cdot\delta_{\hat{y}_{i}-y_{i}}+u^{\prime}\cdot\delta_{\infty}$
4: $L\leftarrow\hat{y}^{\prime}-r$
We adapt conformal prediction techniques to conduct design algorithm selection (Alg. 7). For simplicity of exposition, we will describe the method assuming the function $g$ is the identity; for any other function, one can replace all references to “labels” with $g(Y)$.
We first explain how to construct a valid lower bound for the empirical average of design labels, $(1/N)\sum_{i=1}^{N}y^{\prime}_{i}$, for a single configuration, $\lambda$. We can use a split conformal method [47] to construct valid lower bounds, $l_{i},i=1,\ldots,N$, for the labels of $N$ designs, with confidence of $1-\alpha/N$ each (Alg. 8). These lower bounds each satisfy $\mathbb{P}(y^{\prime}_{i}\geq l_{i})\geq 1-\alpha/N$, where the probability is over random draws of the held-out labeled data. The average of these bounds, $L=(1/N)\sum_{i=1}^{N}l_{i}$, is then a valid lower bound for the empirical average of design labels: $\mathbb{P}((1/N)\sum_{i=1}^{N}y^{\prime}_{i}\geq L)\geq 1-\alpha$. To see why, note that the event $\{y^{\prime}_{i}\geq l_{i},\;\forall i\in[N]\}$ occurs with probability at least $1-\alpha$ due to the Bonferroni correction, and on this event, we have $(1/N)\sum_{i=1}^{N}y^{\prime}_{i}\geq(1/N)\sum_{i=1}^{N}l_{i}$.
To conduct design algorithm selection, we introduce an additional Bonferroni correction for the size of the menu, $|\Lambda|$: for each design from each configuration, we construct a split conformal lower bound with a confidence of $1-\alpha/(|\Lambda|\cdot N|)$. For each configuration, we then take the average lower bound, $L_{\lambda}$. The additional Bonferroni correction means that the event $\{(1/N)\sum_{i=1}^{N}y^{\lambda}_{i}\geq L_{\lambda},\;\forall\lambda\in\Lambda\}$ occurs with probability at least $1-\alpha$, which in turn implies that for $\hat{\Lambda}:=\{\lambda\in\Lambda:L_{\lambda}\geq\tau\}$, we have
@@
Note that the lower bounds, $L_{\lambda},\lambda\in\Lambda$, are immensely conservative: because conformal prediction techniques are tailored explicitly for individual labels, not functionals of label distributions, they cannot naturally account for how signed prediction errors over many designs can “cancel out” in estimation of the mean design label. We rely instead on Bonferroni corrections to guarantee the extremely stringent criterion that each and every individual design’s lower bound is correct. In practice, this meant that $L_{\lambda}$ was always negative infinity in our experiments. Conformal prediction is fundamentally not the right tool when one is interested in how prediction error affects distributions of labels and properties thereof, rather than just the labels of individual examples.
@@ Appendix D Protein GB1 experiment details
D.1 Exceedance-based success criteria
We ran the same experiments described in §5.1, except with success criteria involving the exceedance over $1$: $\mathbb{E}_{Y\sim P_{Y;\lambda}}[\mathbbm{1}(Y\geq 1)]\geq\tau$, for $\tau\in[0,1]$. For context, a label value of $1$ is the $0.994$-quantile of the training labels and represents a binding affinity of about $2.7$ times that of wild type GB1. The results were qualitatively similar to those with success criteria involving the mean design label in the main text (Fig. 6), except that GMMForecasts with $q\in\{0,0.5\}$ was slightly less conservative (though still too conservative to be useful, as it ceased selecting any temperatures for $\tau>0.05$ and $\tau>0.15$ for $q=0$ and $q=0.5$, respectively).
Figure 6: Design algorithm selection for designing protein GB1. (a) Error rate (top; lower is better) and selection rate (bottom; higher is better) for all methods, for range of values of $\tau$, the desired exceedance over $1$ (that is, the fraction of designs whose label exceeds $1$). (b) Median (solid line) and $20^{\text{th}}$ to $80^{\text{th}}$ percentiles (shaded regions), of the lowest mean design label of selected configurations, across trials for which each method did not return the empty set. Dots mark where each median trajectory ends (i.e., the desired exceedance beyond which a method ceases to select any configuration).

### v2

@@ B.1 More General Success Criteria
The main text considers success criteria of the following form: $\theta_{\lambda}:=\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]\geq\tau$ for some $g:\mathcal{Y}\rightarrow\mathbb{R},\tau\in\mathbb{R}$. More generally, we can use prediction-powered inference techniques to compute valid p-values whenever the population-level metric, $\theta_{\lambda}$, can be expressed as the minimizer of the expectation of some convex loss (Angelopoulos et al.,, 2023):
$\theta_{\lambda}:=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[\ell_{\theta}(X,Y)],$
where $\ell_{\theta}$ is convex in $\theta$. When $\ell_{\theta}(X,Y)=(g(Y)-\theta)^{2}$ for some $g:\mathcal{Y}\rightarrow\mathbb{R}$, we recover the special case in the main text. We could not conceive of practical settings requiring this general characterization of $\theta_{\lambda}$, but it may be useful for future work.
@@ B.2 Design and Labeled Densities Known up to Normalizing Constants
We can compute asymptotically valid p-values if we have unnormalized forms of the design and labeled densities, such as when sequences are generated from energy-based models (Biswas et al.,, 2021), Potts models (Russ et al.,, 2020; Fram et al.,, 2024), or other Markov random fields. Specifically, assume we can evaluate $p^{\text{u}}_{X;\lambda}(x)=a\cdot p_{X;\lambda}(x)$ and $p^{\text{u}}_{\text{lab}}(x)=b\cdot p_{\text{lab}}(x)$ for unknown constants $a,b\in\mathbb{R}$, where the superscript indicates that the densities are unnormalized. To leverage these in place of the exact densities in Algorithm 2, consider the corresponding scaled density ratios on the labeled data, $w^{\text{u}}_{j}:=p^{\text{u}}_{X;\lambda}(x_{j})/p^{\text{u}}_{\text{train}}(x_{j}),j=1,\ldots,n$. The self-normalized importance-weighted estimator of prediction bias,
@@
is a consistent estimator of the rectifier in Eq. 4, $\mathbb{E}_{X,Y\sim P_{X;\lambda}\cdot P_{Y\mid X}}[g(Y)-g(f(X))]$. Since $\hat{\Delta}^{\text{u}}$ is a ratio of means, the delta method can be used to derive its asymptotic variance (Owen,, 2013), which can be estimated as
@@
We can then compute an asymptotically valid p-value using Algorithm 2, but replacing $\hat{\Delta}$ with $\hat{\Delta}^{\text{u}}$ in Line 3 and $\hat{\sigma}^{2}_{\text{err}}/n$ with $\hat{\sigma}^{2}_{\hat{\Delta}^{\text{u}}}$ in Line 7.
@@ C.1 Prediction-Only Method
Inputs: Predictions for designs, $\{\hat{y}^{\lambda}_{i}\}_{i=1}^{N}$; desired threshold, $\tau\in\mathbb{R}$.
@@
1: $\hat{\theta}\leftarrow\frac{1}{N}\sum_{i=1}^{N}g(\hat{y}^{\lambda}_{i})$
2: $\hat{\sigma}^{2}_{\text{pred}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}(g(\hat{y}^{\lambda}_{i})-\hat{\theta}))^{2}$
3: $P\leftarrow 1-\Phi\left((\hat{\theta}-\tau)/\sqrt{\hat{\sigma}^{2}_{\text{pred}}/N}\right)$
The prediction-only method runs multiple testing (Alg. 1) with p-values computed using only the predictions for the designs (Alg. 6), treating them as if they were labels. These p-values are asymptotically valid for testing whether $\mathbb{E}_{X\sim P_{X;\lambda}}[g(f(X))]\geq\tau$—that is, whether the expected function of predictions for designs, but not necessarily their labels, surpasses a threshold.
@@ C.2 Gaussian-Mixture-Model Forecasts Method
The GMMForecasts method follows Wheelock et al., (2022), who model the forecast for a designed sequence as a mixture of two Gaussians (representing beliefs over the label if the sequence is “nonfunctional” or “functional”) with sequence-specific mixture proportion, means, and variances. To construct these forecasts, their approach first assumes access to a predictive mean and variance for each designed sequence, as described below. It then uses the training data to fit a mapping from these initial predictive means and variances to Gaussian mixture model (GMM) parameters. It also seeks to address covariate shift between the design and training data by using a sequence’s edit distance from a reference training sequence (set to wild-type GB1 in the GB1 experiments, and the seed sequence in the RNA binder experiments) as an additional feature in fitting this mapping. Forecasts—that is, GMM parameters—are then inferred for each designed sequence.
The forecasts also involve a key hyperparameter, $q\in[0,1]$. For each designed sequence, after GMM parameters are inferred, the mean of the “functional” Gaussian is adjusted by taking a convex combination of it and the original predictive mean, where $q$ weights the latter. Using different values of $q$ reflects how much one trusts the predictions; high values result in forecasts where the “functional” Gaussian mean is determined largely by the original predictive mean. We ran the method with $q\in\{0,0.5,1\}$ to span the range of possible values.
Predictive mean and variance for a sequence
When using predictive models that were ensembles (i.e., the fully connected ensemble in the GB1 experiments, and the fully connected and convolutional ensembles in the RNA binder experiments), the predictive mean and variance for a sequence was set to the empirical mean and variance, respectively, of the predictions for that sequence. When using the ridge regression model in the RNA binder experiments, the model weights were fit on $90\%$ of the training data, and the predictive mean for a sequence was set to the model’s prediction for it. The (homogeneous) predictive variance was set to the model’s mean squared error over the remaining $10\%$ of the training data.
@@ C.3 Calibrated Forecasts Method
The CalibratedForecasts method is based on the idea that forecasts, or models of the conditional distributions of the label, should be statistically consistent with held-out labeled data. Specifically, we aim for forecasts to be calibrated as defined by Kuleshov et al., (2018), whose definition is related to the notion of probabilistic calibration (Gneiting et al.,, 2007):
@@
where $F_{X}$ denotes the CDF of the forecast for a sequence $X\in\mathcal{X}$, and the probability is over the distribution of labeled data. That is, for any $p\in[0,1]$, the label falls under the $p$-quantile given by a calibrated forecast with frequency $p$.
For a given sequence, let an initial forecast be a Gaussian with mean and variance set to a predictive mean and variance, as described above for GMMForecasts . We then use the held-out labeled data to learn a transformation of these initial forecast CDFs, such that the transformed CDF is closer to achieving calibration (Eq. 12). Specifically, we use isotonic regression, following Kuleshov et al., (2018). We then construct a forecast for every designed sequence by first forming the initial forecast, and then transforming the corresponding CDF with the fitted isotonic regression function.
@@ C.4 Conformal Prediction Method
Inputs: Designs generated with each configuration, $\{x^{\lambda}_{i}\}_{i=1}^{N}$ for all $\lambda\in\Lambda$; predictive models used by each configuration, $\{f_{\lambda}\}_{\lambda\in\Lambda}$; held-out labeled data, $\{(x_{j},y_{j})\}_{j=1}^{n}$; desired threshold, $\tau\in\mathbb{R}$; error rate, $\alpha\in[0,1]$.
@@
2:      Predictions for designs, $\hat{y}^{\lambda}_{i}\leftarrow f_{\lambda}(x^{\lambda}_{i}),i=1,\ldots,N$
3:      Predictions for labeled data, $\hat{y}_{j}\leftarrow f_{\lambda}(x_{i}),j=1,\ldots,n$
4:      Density ratios for designs, $v_{i}\leftarrow\textsc{DensityRatio}_{\lambda}(x^{\lambda}_{i}),i=1,\ldots,N$
5:      Density ratios for labeled data, $w_{j}\leftarrow\textsc{DensityRatio}_{\lambda}(x_{j}),j=1,\ldots,n$
6:      $l^{\lambda}_{i}\leftarrow\textsc{SplitConformalLB}((\hat{y}^{\lambda}_{i},v_{i}),\;\{(y_{j},\hat{y}_{j},w_{j})\}_{j=1}^{n},\;\alpha/(|\Lambda|\cdot N)),i=1,\ldots,N$
7:      $L_{\lambda}\leftarrow(1/N)\sum_{i=1}^{N}l^{\lambda}_{i}$
@@
Inputs: a design’s prediction and density ratio, $(\hat{y}^{\lambda},v)$; labels, predictions, and density ratios for held-out labeled data, $\{(y_{j},\hat{y}_{j},w_{j})\}_{j=1}^{n}$; $\alpha\in[0,1]$.
@@
1: $u_{j}\leftarrow\frac{w_{j}}{\sum_{j=1}^{n}w_{j}+v},j=1,\ldots,n$
2: $u\leftarrow\frac{v}{\sum_{j=1}^{n}w_{j}+v}$
3: $r\leftarrow\text{$(1-\alpha)$-quantile of the distribution comprising the mixture of point masses}\sum_{j=1}^{n}u_{j}\cdot\delta_{\hat{y}_{j}-y_{j}}+u\cdot\delta_{\infty}$
4: $L\leftarrow\hat{y}^{\lambda}-r$
We adapt conformal prediction techniques to conduct design algorithm selection (Algs. 7,  8). For simplicity, we describe the method with $g$ as the identity function; for other functions, one can replace all references to labels with $g(Y)$.
For a given configuration, $\lambda$, we can construct a valid lower bound for the empirical average of design labels, $(1/N)\sum_{i=1}^{N}y^{\lambda}_{i}$, by averaging Bonferroni-corrected conformal lower bounds for each design. Concretely, we can use a split conformal method (Tibshirani et al.,, 2019) to construct lower bounds, $l_{i}$, for the labels of $N$ designs, with confidence of $1-\alpha/N$ each (Alg. 8). These lower bounds each satisfy $\mathbb{P}(y^{\lambda}_{i}\geq l_{i})\geq 1-\alpha/N$, where the probability is over random draws of the held-out labeled data. The average of these bounds, $L=(1/N)\sum_{i=1}^{N}l_{i}$, then satisfies $\mathbb{P}((1/N)\sum_{i=1}^{N}y^{\lambda}_{i}\geq L)\geq 1-\alpha$. This is because the event $\{y^{\lambda}_{i}\geq l_{i},\;\forall i\in[N]\}$ occurs with probability at least $1-\alpha$ due to the Bonferroni correction, and on this event, we have $(1/N)\sum_{i=1}^{N}y^{\lambda}_{i}\geq(1/N)\sum_{i=1}^{N}l_{i}$.
To conduct design algorithm selection, we introduce an additional Bonferroni correction for the size of the menu, $|\Lambda|$: for each configuration, we construct a split conformal lower bound with a confidence of $1-\alpha/(|\Lambda|\cdot N|)$ for each design, then take their average lower bound, $L_{\lambda}$. The event $\{(1/N)\sum_{i=1}^{N}y^{\lambda}_{i}\geq L_{\lambda},\;\forall\lambda\in\Lambda\}$ occurs with probability at least $1-\alpha$, which in turn implies that for $\hat{\Lambda}:=\{\lambda\in\Lambda:L_{\lambda}\geq\tau\}$, we have
@@
Note that the lower bounds $L_{\lambda},\lambda\in\Lambda$, are immensely conservative: because conformal prediction techniques are meant to characterize individual labels, they cannot naturally account for how prediction errors over many designs can “cancel out” in estimation of the mean design label. We rely instead on Bonferroni corrections to guarantee the extremely stringent criterion that every individual design’s lower bound is correct, which in practice meant that $L_{\lambda}$ was always negative infinity in our experiments. Conformal prediction is fundamentally not the right tool when one is interested in how prediction error affects distributions of labels, rather than the labels of individual instances.
@@ Appendix D Protein GB1 Experiment Details
D.1 Estimated Density Ratios
We ran our method with estimated density ratios, for success criteria concerning the mean design label. Specifically, we separately estimated the labeled distribution and the design distribution corresponding to each configuration. For the former, we performed maximum-likelihood estimation with Laplace smoothing (with pseudocounts of 1) to estimate the site-specific categorical distributions using the held-out labeled sequences; for the latter, we did the same using the designed sequences. For a given sequence, we then took the ratio of its densities under these two estimated distributions as the estimated density ratio. The results from our method using these estimated density ratios were very similar to the original results using the known density ratios (Fig. 6).
Figure 6: Design algorithm selection for designing protein GB1 (estimated density ratios). (a) Error rate (top; lower is better) and selection rate (bottom; higher is better) for our original results (with known density ratios between the designed and labeled distributions; green) and our method with estimated density ratios (purple). (b) The median (solid line) and $20^{\text{th}}$ to $80^{\text{th}}$ percentiles (shaded regions) of the lowest mean design label achieved by selected configurations, over trials for which the method did not return the empty set. Dots mark where each median trajectory ends (i.e., the value of $\tau$ beyond which the method ceased to select any configuration, and the lowest mean design label of configurations selected for that $\tau$). Results on or above the dashed diagonal line indicate that selected configurations are successful.
D.2 Exceedance-Based Success Criteria
We ran the same experiments described in §5.1 and shown in Figure 3, except with success criteria involving the exceedance over $1$: $\mathbb{E}_{Y\sim P_{Y;\lambda}}[\mathbbm{1}(Y\geq 1)]\geq\tau$, for $\tau\in[0,1]$. For context, a label value of $1$ was greater than $99.4\%$ of the training labels and represents a binding affinity of about $2.7$ times that of wild-type GB1. The results were qualitatively similar to those in the main text, except that GMMForecasts with $q\in\{0,0.5\}$ was slightly less conservative and yielded high selection rates for a limited range of $\tau$ (Fig. 7).
Figure 7: Design algorithm selection for designing protein GB1 (exceedance-based success criteria). (a) Error rate (top; lower is better) and selection rate (bottom; higher is better) for all methods, for range of values of $\tau$, the desired exceedance over $1$ (that is, the fraction of designs whose label exceeds $1$). For reference, the training labels’ exceedance over $1$ was $0.006$. GMMForecasts with hyperparameter $q\in\{0,0.5,1\}$ (dark purple, medium purple, and pink lines) had error rates of zero for all values of $\tau$. (b) Median (solid line) and $20^{\text{th}}$ to $80^{\text{th}}$ percentiles (shaded regions), of the lowest mean design label achieved by selected configurations, across trials for which each method did not return the empty set. Dots mark where each median trajectory ends (i.e., the value of $\tau$ beyond which a method ceases to select any configuration, and the lowest mean design label of configurations selected for that $\tau$). Results on or above the dashed diagonal line indicate that selected configurations are successful.